### v1

# A non-autonomous Hamiltonian diffeomorphism with roots of all orders

Abstract.
We provide an example of a non-autonomous Hamiltonian diffeomorphism with roots of all orders. This answers a question of Egor Shelukhin.
1. Introduction
1.1. Main result and motivation
We say that a Hamiltonian diffeomorphism is autonomous when it is the time-1 map of a flow generated by an autonomous Hamiltonian, i.e. time-independent. The set of all autonomous Hamiltonian diffeomorphisms is denoted by $\text{Auto}(M,\omega)$. Autonomous Hamiltonian diffeomorphisms enjoy properties that are generally not true for general Hamiltonian diffeomorphisms; for example, one can show that if $H$ is an autonomous Hamiltonian, then each level set $H^{-1}(\{c\})$ for $c\in\mathbb{R}$ is preserved under the flow $\varphi^{t}_{H}$.
In many ways, the set $\text{Auto}(M,\omega)$ is believed to be small in the group of all Hamiltonian diffeomorphisms, $\text{Ham}(M,\omega)$, as pointed out in [PS16], this is a symplectic take on Palis’ heuristic [Pal74]: “Vector fields generate few diffeomorphisms”. Polterovich and Shelukhin [PS16], followed by Polterovich, Shelukhin and Stojisavljević [PSS17], Zhang [Zha19], Chor [Cho20], and Khanevsky [Kha22] indeed showed that for a wide class of closed symplectic manifolds, there exist Hamiltonian diffeomorphisms arbitrarily far from elements of $\text{Auto}(M,\omega)$ with respect to the Hofer norm. A key observation for the proof of the result by Polterovich and Shelukhin [PS16] is that all autonomous Hamiltonian diffeomorphisms admit roots of all orders. This comes from the fact that if the time-1 map of the Hamiltonian flow generated by $H\colon M\to\mathbb{R}$ is $\varphi$, then the time-1 map of the Hamiltonian flow generated by $H/k$ is a $k$-th root of $\varphi$. It is thus a natural question to ask whether the only maps that has roots of all order are autonomous. Shelukhin has asked this question in [She24, Section 3] even though the question has been known in the community since around 2015 [Sto]. In this paper, we show the existence of non-autonomous Hamiltonian diffeomorphisms of general symplectic manifolds that admit roots of all orders. Our main result is the following.
Theorem 1.
For any symplectic manifold $(M,\omega)$ there exists a group homomorphism
$\Phi\colon\mathbb{Q}\to\text{Ham}(M,\omega),$
such that
$\Phi(\mathbb{Q})\cap\text{Auto}(M,\omega)=\{Id\}.$
For $\varphi:=\Phi(1)$ the group homomorphisms property of $\Phi$ directly implies that
$\left(\Phi\left(1/k\right)\right)^{k}=\Phi(1)=\varphi.$
In other words, $\Phi(1/k)$ is a $k$-th root of $\varphi$. This implies the following corollary that answers Shelukhin’s question.
Corollary 1.1.
For any symplectic manifold $(M,\omega)$, there exists a Hamiltonian diffeomorphism $\varphi$ that has roots of all orders but is not autonomous.
We will see that the image of the constructed group homomorphisms $\Phi$ lies in the $C^{\infty}$-closure of the set of autonomous diffeomorphisms. Naturally, we ask the following question.
Question 1.
Does there exist a Hamiltonian diffeomorphism that has roots of all orders but does not lie in the $C^{\infty}$-closure of the set of autonomous diffeomorphisms?
As we will show shortly, the proof of Theorem 1 is a consequence of the following proposition.
Proposition 1.2.
For any symplectic manifold $(M,\omega)$ there exists a group homomorphism
$\Phi\colon\mathbb{Q}\to\text{Ham}(M,\omega)$
and an open set $\mathcal{U}\subset M$ such that for all $\ell\in\mathbb{Z}_{\geq 1}$ there exists a point $x\in M$ with a positive orbit
$\mathcal{O}(\Phi(\ell),x):=\{x,\Phi(\ell)(x),\Phi(2\ell)(x),\ldots\}$
that is dense in $\mathcal{U}$.
This proposition implies Theorem 1, as we show now. The existence of orbits of $\Phi(\ell)$ that are dense in an open set implies that they are not autonomous.
Proof of Theorem 1.
Suppose that there exists an autonomous Hamiltonian diffeomorphism $\varphi\in\text{Auto}(M,\omega)$, generated by $H$ such that there exists a point $x$ of $M$ with an orbit $\mathcal{O}(\varphi,x)$ that is dense in an open set $\mathcal{U}\subset M$. Since autonomous Hamiltonians are preserved by the action of their flow, we have that $H$ is constant on $\mathcal{O}(\varphi,x)$. However, since the orbit is dense in $U$, the Hamiltonian $H|_{U}$ restricted to $U$ must be constant. Consequently, $\varphi$ acts as the identity on $U$, and the orbit $\mathcal{O}(\varphi,x)$ consist of only one point. This contradicts the density of the orbit in the open set $U$.
After applying this argument to $\Phi(\ell)$ for any $\ell\in\mathbb{Z}_{\geq 1}$ we can conclude that
$\Phi(\mathbb{Q})\cap\text{Auto}(M,\omega)=\{Id\}$
and therefore finish the proof of Theorem 1 assuming Proposition 1.2. ∎
1.2. Organization of the paper
The paper is organized as follows. In Section 2, we prove Proposition 1.2 by adapting the Anosov-Katok construction to our setting. In Section 3 we study the continuity of the map defined in Theorem 1 for natural topologies. In this last section we provide only partial answers.
1.3. Acknowledgments
The authors would like to thank P. Feller, F. Le Roux, L. Polterovich, V. Stojisavljević and I. Trifa for their interest and comments about this work. The authors would like to thank especially E. Çinelli and S. Seyfaddini for comments and insightful discussion throughout the project. The second author would also like to thank M. Burger for a discussion that led him to understand better the mathematical context of this paper as well as A. Ulliana and H. Wu for organizing a seminar that contributed to the birth of this paper. During this project, B.S. was supported by ERC Starting Grant 851701.
2. Proof of Proposition 1.2
This section is dedicated to the proof of Proposition 1.2. We recall first some basic definitions and notations of symplectic geometry. We then describe our set-up for the Anosov-Katok construction that differs from the usual one. We finish the proof of Proposition 1.2 in Section 2.3 and 2.4.
2.1. Preliminaries of symplectic geometry
A symplectic manifold $M$ is a manifold endowed with a symplectic form $\omega$, i.e. a closed non-degenerate 2-form. For any symplectic manifold $(M,\omega)$, we denote by $\text{Symp}(M,\omega)$ the group of symplectomorphisms of $M$. Let $H_{t}\colon M\to\mathbb{R}$ be a time-dependent compactly supported Hamiltonian; one defines the Hamiltonian vector field $X_{H}$ associated to $H$ to be the unique vector field that verifies the following formula
$\iota_{X_{H}}\omega=-\mathrm{d}H.$
The flow of the vector field $X_{H}$ is called the Hamiltonian flow of $H$ and is denoted $\varphi^{t}_{H}$. The time-1 maps of Hamiltonian flows are called Hamiltonian diffeomorphisms, and altogether, they form a subgroup of $\text{Symp}(M,\omega)$ denoted by $\text{Ham}(M,\omega)$.
2.2. Set-up of the Anosov-Katok construction
The proof consists of constructing a Hamiltonian diffeomorphism with the Anosov-Katok construction, introduced first in [AK70]. Their construction uses a Hamiltonian $S^{1}$-action on a symplectic manifold. We generalize the construction to perform it on symplectic manifolds $(M,\omega)$ with a smooth Hamiltonian action $S:\mathbb{R}\to(\text{Ham}(M,\omega),C^{\infty})$ such that there exists an invariant open subset $U\subset M$ upon which the restriction of the Hamiltonian action $S(\cdot)|_{U}$ is 1-periodic and induces a free circle action without fixed points denoted by
$S|_{U}:S^{1}\to\text{Symp}(U,\omega).$
With this data, we will construct a sequence of group homomorphisms $(\Phi_{n}:\mathbb{R}\to\text{Ham}(M,\omega))_{n}$ such that the following hold.
  • (i)
    The group homomorphisms $\Phi_{n}$ are autonomous flows.
  • (ii)
    For any $q\in\mathbb{Q}$ the sequence $(\Phi_{n}(q))_{n}$ converges in the $C^{\infty}$-topology. The limits of all rational points fit together to form a group homomorphism
    $\Phi:\mathbb{Q}\to\text{Ham}(M,\omega).$
  • (iii)
    For every $\ell\in\mathbb{Z}_{\geq 1}$, the diffeomorphisms $\Phi(\ell)$ have a dense orbit in $U$.
The existence of such a group homomorphism $\Phi:\mathbb{Q}\to\text{Ham}(M,\omega)$ proves Proposition 1.2. In the subsequent two sections, we first show how the sequence $(\Phi_{n})_{n}$ that satisfies (i) and (ii) is constructed and then argue that we can improve the construction, additionally achieving (iii).
Let us assume for now that the above construction can be carried out, then once a suitable group action $S\colon\mathbb{R}\to\text{Ham}(M,\omega)$ exists, we have proven Proposition 1.2. Such action exists by definition on every symplectic manifold that carries a non-trivial Hamiltonian circle action. However, there exist symplectic manifolds that do not admit a non-trivial Hamiltonian $S^{1}$-action. We now show how to construct a suitable action $S$ on every symplectic manifold.
Our construction will take place in $\mathbb{R}^{2n}$ endowed with the standard symplectic form $\omega_{0}$. Darboux’s theorem allows us to embed this construction in any symplectic manifold. Let $u_{0}:\mathbb{R}_{\geq 0}\to[0,1]$ be a function given by
$u_{0}(x)=\begin{cases}1-x,&\text{for }x\in[0,1]\\ 0,&\text{otherwise. }\end{cases}$
We now locally smoothen $u_{0}$ in $[0,\varepsilon)$ and $(1-\varepsilon,1+\varepsilon)$, two small neighborhoods of 00 and $1$ respectively. By doing this, we can obtain a smooth compactly supported function $u:\mathbb{R}_{\geq 0}\to[0,1]$ that is locally constant at 00 and coincides with $u_{0}$ on $(\varepsilon,1-\varepsilon)$. We now define the radial Hamiltonian on $\mathbb{R}^{2n}$ by
$G(x)=u(|x|).$
The flow generated by $G$ is a suitable Hamiltonian action $S:\mathbb{R}\to\text{Ham}(\mathbb{R}^{2n},\omega_{0})$. Namely, denoting $B_{r}$ for the open ball of radius $r$ centered at the origin, the open set $U=B_{1-\varepsilon}-\overline{B_{\varepsilon}}$ is invariant under $S$ and $S$ acts on $U$ as a free circle action.
2.3. Scheme of the construction
We construct the sequence of autonomous morphisms $\Phi_{n}$ with the Anosov-Katok method, following the scheme described in [RS22]. Our construction differs from theirs in that we use a diagonal argument to impose the convergence of a countable number of maps simultaneously. We denote by $d_{C^{\infty}}$ a distance on $\text{Symp}(\text{supp}(S),\omega)$ that induces the Whitney $C^{\infty}$-topology and turns it into a complete metric space.
The backbone of this construction is the Hamiltonian action $S$ that restricts to a circle action on an invariant set $U$. Then the morphisms $\Phi_{n}$ are defined as conjugations of the action $S$ stretched by the factor $\alpha_{n}$. In other words, we have
$\Phi_{n}(\lambda)=H_{n}S_{\lambda\alpha_{n+1}}H_{n}^{-1},$
where $H_{n}\in\text{Ham}(U,\omega)$ is a Hamiltonian diffeomorphism supported in $U$ and $\alpha_{n}\in\mathbb{Q}$. On top of this, we also demand that when written in irreducible form $\alpha_{n}=\frac{p_{n}}{q_{n}}$ is such that $p_{n}$ is divisible by $n!$. In what follows we will need fast convergence of the sequence of rational numbers $\alpha_{n}$, we describe briefly why the constraint on the numerator is harmless with respect to that.
Lemma 2.1.
Let $n$ be any integer, then the set of fractions that can be written with a numerator that is divisible by $n!$ in their irreducible representation is dense in $\mathbb{R}$.
Proof.
Let $\varepsilon>0$ be an positive number, pick $q$ a prime number that is bigger than $n!$ and $\frac{2n!}{\varepsilon}$. Then, the set of all irreducible fractions with denominator $q$ and numerator divisible by $n!$ forms is an $\varepsilon$-net inside $\mathbb{R}$. Since we can do this for all $\varepsilon>0$, the result follows. ∎
The construction will be inductive; we construct the Hamiltonian diffeomorphism $H_{n}$ such that $H_{n}:=H_{n-1}\circ h_{n}$ for some Hamiltonian diffeomorphism $h_{n}\in\text{Ham}(U,\omega)$ that commutes with $S_{1/q_{n}}$. This way for every $\lambda\in\mathbb{Q}$ with denominator in $\llbracket 1,n\rrbracket=[1,n]\cap\mathbb{Z}$, the following holds
(1) $\Phi_{n-1}(\lambda)=H_{n-1}S_{\lambda\alpha_{n}}H_{n-1}^{-1}=H_{n}S_{\lambda\alpha_{n}}H_{n}^{-1}.$
The identity (1) implies that if we choose $\alpha_{n+1}$ close enough to $\alpha_{n}$, we can conclude that for all $k\in\llbracket 1,n\rrbracket$, we have
$d_{C^{\infty}}\left(\Phi_{n}\left(1/k\right),\Phi_{n-1}\left(1/k\right)\right)$ $=d_{C^{\infty}}(H_{n}S_{\alpha_{n+1}/k}H_{n}^{-1},H_{n-1}S_{\alpha_{n}/k}H_{n-1}^{-1})$
$=d_{C^{\infty}}(H_{n}S_{\alpha_{n+1}/k}H_{n}^{-1},H_{n}S_{\alpha_{n}/k}H_{n}^{-1})$
(2) $\leq\dfrac{1}{2^{n}}.$
Note that from the imposed condition on the numerator of $\alpha_{n}$ it follows that for all $\lambda\in\mathbb{Q}$ the equation (2) will hold eventually for large enough $n$. This means that that for every $\ell\in\mathbb{Z}_{\geq 1}$, the sequence of maps $(\Phi_{n}({1\over\ell}))_{n}$ is a Cauchy sequence and thus converges in $C^{\infty}$-topology to a diffeomorphism that we denote $\Phi({1\over\ell})$. Given the pointwise convergence of the sequence $(\Phi_{n})_{n}$ on $\mathbb{Q}$ and the fact that each $\Phi_{n}$ is a group homomorphism, it follows directly that the sequence converges to a map $\Phi\colon\mathbb{Q}\to\text{Diff}(M,\omega)$, which is also a group homomorphism. Since the $C^{k}$-flux conjecture holds for $1\leq k\leq\infty$ [Ono06], we can conclude that the limit diffeomorphisms lie in $\text{Ham}(M,\omega)$, that is we have constructed
$\Phi:\mathbb{Q}\to\text{Ham}(M,\omega).$
This construction is very general; one can now demand additional properties for the sequences $(\alpha_{n})_{n}$ and $(H_{n})_{n}$ to achieve specific properties in the limit $\Phi$. For example, as described in [RS22], by demanding light conditions for $H_{n}$, the limit diffeomorphisms can be forced to have a dense orbit in $U$ and, therefore, are not autonomous.
Although the condition on $h_{n}$ might seem very strong, it can be reformulated as follows. For any $n\in\mathbb{Z}_{\geq 1}$, we can consider the quotient $(U/S_{1/q_{n}},\sigma)$ which is a smooth symplectic manifold with symplectic structure $\sigma$ induced by $\omega$. Note that there is a smooth covering
$\pi:U\to U/S_{1/q_{n}}.$
Any Hamiltonian diffeomorphism $g_{n}\in\text{Ham}(U/S_{1/q_{n}},\sigma)$ admits a lift $h_{n}\in\text{Ham}(U,\omega)$ that commutes with $S_{1/q_{n}}$ and therefore (1) is achieved.
2.4. Achieving dense orbits
The proof that a dense orbit in the limit diffeomorphisms of the Anosov-Katok construction can be achieved as described in [RS22]. For the convenience of the reader and since we want the limit diffeomorphisms $\Phi(\ell)$ to have dense orbits for all $\ell\in\mathbb{Z}_{\geq 1}$, we describe how this construction can be achieved.
Fix an auxiliary Riemannian metric on $M$ and denote the induced distance by $d$. A subset $F\subset U$ is called $\varepsilon$-dense in $U$ if
$\sup_{x\in U}\inf_{y\in F}d(x,y)<\varepsilon.$
Now choose a fixed sequence $(\varepsilon_{n})_{n}$ of real numbers converging to 00. We will choose $g_{n}\in\text{Ham}(U/S_{1/q_{n}},\sigma)$ such that $\Phi_{n}(n!)$ has an orbit $\mathcal{O}(\Phi_{n}(n!),x_{n})$ that is $\varepsilon_{n}$-dense in $U$. Note that since for $\ell=1,...,n$ we have
$\mathcal{O}(\Phi_{n}(n!),x_{n})\subset\mathcal{O}(\Phi_{n}(\ell),x_{n}),$
and thus $\Phi_{n}(\ell)$ has an $\varepsilon_{n}$-dense orbit in $U$. Fix $\ell\in\mathbb{Z}_{\geq 1}$, for $n$ an integer large enough, the map $\Phi_{n}(\ell)$ has an $\varepsilon_{n}$-dense orbit, we obtain as in [RS22, Section 2.3] that fast convergence implies that $\Phi(\ell)$ has an $\varepsilon$-dense orbit in $U$ for any $\varepsilon>0$. Again as in [RS22] it implies that $\Phi(\ell)$ has a dense orbit in $U$. This would imply Proposition 1.2.
We still need to show that the sequence $(g_{n})_{n}$ can be chosen such that $\Phi_{n}(n!)$ has an $\varepsilon_{n}$-dense orbit. We assume that we already constructed $\alpha_{n}$ and $H_{n-1}$ such that the map $\Phi_{n-1}((n-1)!)=H_{n-1}S_{\alpha_{n}(n-1)!}H_{n-1}^{-1}$ has an $\varepsilon_{n-1}$-dense orbit and construct $H_{n}$ and $\alpha_{n+1}$ such that $\Phi_{n}(n!)$ has an $\varepsilon_{n}$-dense orbit. Set
$\pi_{n}:U\to U^{\prime}:=U/S_{1/q_{n}}$
to be the standard projection. Choose a finite set of points $F_{1}\subset U^{\prime}$ such that $\pi_{n}^{-1}(F_{1})$ is $\eta_{n}$-dense in $U$ where $\eta_{n}$ is chosen such that $H_{n-1}$ maps any ball of diameter $\eta_{n}$ inside a ball of diameter $\varepsilon_{n}$. Note that this is possible since $U$ lies in the support of the flow $S$, which is compact. We choose a set $F_{1}\subset U^{\prime}$ with the same cardinality as $F_{1}$, all lying in the projection $\pi(C)$ of a single orbit $C$ of the circle action on $U$. Since the group of Hamiltonian diffeomorphisms is $|F_{1}|$-transitive (Lemma 2.2), we can pick $g_{n}\in\text{Ham}(U^{\prime})$ such that $g_{n}(F_{1})=F_{2}$.
Lemma 2.2.
For a symplectic manifold $(M,\omega)$ and any $p$-tuples $(x_{1},...,x_{p})$ and $(y_{1},...,y_{p})$ where $x_{i}\neq x_{j}$ and $y_{i}\neq y_{j}$ for $i\neq j$ there exists a compactly supported Hamiltonian diffeomorphism $\psi$ such that $\psi(x_{i})=y_{i}$ for all $i$.
We now define $h_{n}$ as the lift of $g_{n}$ to $U$. The image $h_{n}(C)$ of the orbit $C$, is $\eta_{n}$-dense in $U$ and thus, $H_{n-1}(h_{n}(C))$ is $\varepsilon_{n}$-dense in $U$. We choose a rational number $\alpha_{n+1}=\frac{p_{n+1}}{q_{n+1}}$ with $p_{n+1}$ divisible by $(n+1)!$ such that there is a discrete orbit $C^{\prime}\subset C$ under $S_{\alpha_{n+1}n!}$ that contains a subset close enough to $\pi^{-1}(F_{1})$ such that $H_{n-1}(h_{n}(C^{\prime}))$ is $\varepsilon_{n}$-dense in $U$, this is possible if $\alpha_{n+1}$ is closed enough to $\alpha_{n}$. We define $H_{n}:=H_{n-1}\circ h_{n}$ and $\Phi_{n}=H_{n}S_{\alpha_{n+1}}H_{n}^{-1}$. The set $H_{n}(C^{\prime})$ is then an $\varepsilon_{n}$-dense orbit of $\Phi_{n}(n!)$. This concludes the proof of Proposition 1.2.
3. Relation with the $\gamma$-topology and Hofer-topology
We attempted to answer the following question by studying the group homomorphisms $\Phi$ as in Theorem 1.
Question 2.
Is the map $\Phi\colon\mathbb{Q}\to\text{Ham}(M,\omega)$ constructed in Theorem 1 $C^{0}$-continuous? Is it $C^{\infty}$-continuous?
If the map $\Phi$ were to be continuous for the $C^{k}$-topology where $0\leq k\leq\infty$, then it extends to a continuous group morphism from the real line to some completion of $\text{Ham}(M,\omega)$. This map would be nowhere differentiable; otherwise, it would be given by the flow of an autonomous Hamiltonian vector field. Hence, the extension of $\Phi$ would be an example of a morphism from the real line to a completion of $\text{Ham}(M,\omega)$ that is continuous but nowhere differentiable. Unfortunately, we were not able to say much about the relation of $\Phi$ with any $C^{k}$-topologies on $\text{Ham}(M,\omega)$. However, the question becomes more manageable when we ask for the Hofer- or $\gamma$-topology on $\text{Ham}(\mathbb{R}^{2n},\omega_{0})$. We recall some properties of the $\gamma$-norm and the Hofer norm. The Hofer norm is defined in the following way, for all Hamiltonian diffeomorphisms $\varphi\in\text{Ham}(M,\omega),$
$d_{H}(Id,\varphi):=\inf\left\{\int_{0}^{1}(\max_{M}H_{t}-\min_{M}H_{t})\mathrm{d}t\mid H_{t}\text{ generates a path }\varphi_{t},~{}\varphi_{0}=Id,~{}\varphi_{1}=\varphi\right\}.$
As the notation already indicates, the Hofer norm is induced by a metric $d_{H}$ called the Hofer-metric. One can see that the Hofer metric is indeed a metric and it is bi-invariant. The fact that it is non-degenerate is a deep result proven by Lalonde and McDuff in [LM95].
The $\gamma$-norm was first defined by Viterbo in [Vit92] using the theory of generating functions for Lagrangians in the cotangent bundle of a manifold, and it satisfies the following properties.
Proposition 3.1 (Viterbo).
There exists a map $\gamma\colon\text{Ham}(\mathbb{R}^{2n},\omega_{0})\to\mathbb{R}$ that verifies the following properties for all $\varphi,\psi\in\text{Ham}(\mathbb{R}^{2n})$.
  • (i)
    $\gamma(\varphi\psi)\leq\gamma(\varphi)+\gamma(\psi)$.
  • (ii)
    $\gamma(\varphi)\geq 0$ and $\gamma(\varphi)=0$ if and only if $\varphi=Id$.
  • (iii)
    $\gamma(\psi\varphi\psi^{-1})=\gamma(\varphi)$.
  • (iv)
    $\gamma(\varphi)\leq d_{H}(\varphi,Id)$.
  • (v)
    If $H\leq K$ are two Hamiltonians, then $\gamma(\varphi^{1}_{H})\leq\gamma(\varphi^{1}_{K})$.
The normed space $(\text{Ham}(\mathbb{R}^{2n},\omega_{0}),\gamma)$ is not complete and we denote $\widehat{\text{Ham}}(\mathbb{R}^{2n},\omega_{0})$ its completion called the Humilière-completion. Note that elements of $\widehat{\text{Ham}}(\mathbb{R}^{2n},\omega_{0})$ need not be maps. It turns out that $\Phi$ as in Theorem 1 constructed above is both $\gamma$- and Hofer-continuous and hence can be extended over the completion, this is the content of the next proposition. Proposition 3.2 and Proposition 3.3 are formulated for the $\gamma$-norm, but they can also be formulated in terms of the Hofer norm. This choice was made because the completion of $\text{Ham}(M,\omega)$ for the Hofer-norm is not studied as much as the Humilière-completion.
Proposition 3.2.
The map
$\Phi\colon\mathbb{Q}\to\text{Ham}(\mathbb{R}^{2n},\omega_{0})$
as constructed above is $\gamma$-continuous. As a consequence, $\Phi$ extends as a homomorphism in the Humilière-completion
$\widehat{\Phi}\colon\mathbb{R}\to\widehat{\text{Ham}}(\mathbb{R}^{2n},\omega_{0}).$
Moreover, the map $\widehat{\Phi}$ is injective.
Proof.
Since $\Phi$ is a group homomorphism, it is sufficient to show that $\Phi$ is $\gamma$-continuous at 0.
Let $\Phi_{n}(t)=H_{n}S_{\alpha_{n}t}H_{n}^{-1}$ be the sequence of map constructed in Section 2 and let $r$ be a rational number. By construction, we know that the sequence of maps $\Phi_{n}(r)$ converges in the $C^{\infty}$-topology to $\Phi(r)$, hence also in the $\gamma$-topology. This and the conjugation-invariance of the $\gamma$-norm implies that
$\gamma(\Phi(r))$ $=\lim_{n\to\infty}\gamma(\Phi_{n}(r))$
$=\lim_{n\to\infty}\gamma(H_{n}S_{\alpha_{n}r}H_{n}^{-1})$
$=\lim_{n\to\infty}\gamma(S_{\alpha_{n}r})$
$=\gamma(S_{\alpha r})$
where $\alpha\in\mathbb{R}$ is the limit of the sequence $(\alpha_{n})_{n}$. Therefore, we find that $\lim_{r\to\infty}\gamma(\Phi(r))=0$ as expected. Thus, we can define $\widehat{\Phi}\colon\mathbb{R}\to\widehat{\text{Ham}}(\mathbb{R}^{2n},\omega_{0})$ the extension of $\Phi$ to the real line. To show that the map is injective, we need to notice that for all $t\in\mathbb{R}$, $\gamma(\widehat{\Phi}(t))=\gamma(S_{\alpha t})$ vanishes if and only if $S_{\alpha t}=Id$ that is if and only if $t=0$.
Proposition 3.3.
The image of the extension $\widehat{\Phi}$ of Proposition 3.2 is closed for the $\gamma$-topology inside $\widehat{\text{Ham}}(\mathbb{R}^{2n},\omega_{0})$.
Proof.
Let $\varphi\in\widehat{\text{Ham}}(\mathbb{R}^{2n},\omega_{0})$ be a point in the $\gamma$-closure of $\text{Im}(\widehat{\Phi})$. We show that $\varphi$ is in $\text{Im}(\widehat{\Phi})$. Since $\widehat{\text{Ham}(\mathbb{R}^{2n},\omega)}$ is a metric space, there exists a sequence $x_{n}\in\mathbb{R}$ such that the sequence $(\widehat{\Phi}(x_{n}))_{n}$ $\gamma$-converges to $\varphi$. Hence, the sequence $(\widehat{\Phi}(x_{n}))_{n}$ is a Cauchy sequence and therefore
[…]
Using the fact that we have a group homomorphism, we get
[…]
However, as shown above, $\gamma(\widehat{\Phi}(x_{n}-x_{m}))=\gamma(S_{\alpha(x_{n}-x_{m})})$. Since $S$ is generated by a positive Hamiltonian in our construction (see the beginning of Section 2), the map $t\mapsto\gamma(S_{\alpha t})$ is increasing on $\mathbb{R}_{\geq 0}$. This implies that
$x_{n}-x_{m}\xrightarrow[n,m\to+\infty]{}0,$
in other words the sequence $(x_{n})_{n}$ is a Cauchy sequence itself. In particular, $(x_{n})_{n}$ converges to $x\in\mathbb{R}$ and by $\gamma$-continuity,
[…]
Thus, $\varphi$ is indeed in $\text{Im}(\widehat{\Phi})$. ∎
We ask the following natural question that is weaker than Question 2.
Question 3.
Can we associate elements of $\text{Im}(\widehat{\Phi})$ with honest maps? Are they homeomorphisms? Are they diffeomorphisms?
3.1. The special case of manifolds with a Hamiltonian circle action
In this section, we restrict our attention to the case where the symplectic manifold $(M,\omega)$ admits a Hamiltonian circle action $S\colon S^{1}\to\text{Ham}(M,\omega)$. In this setting, we can adapt Proposition 3.1 and Proposition 3.2 to prove the following. Note that the $\gamma$-norm can be defined on every symplectic manifold using Floer Homology by a work of Fukaya and Ono, [FO99]. In general, the $\gamma$-norm enjoys the same properties as the one defined by Viterbo except for property (v) of Proposition 3.1.
Proposition 3.4.
Let $(M,\omega)$ be a symplectic manifold with a Hamiltonian circle action $S$. Then, the map
$\Phi\colon\mathbb{Q}\to\text{Ham}(M,\omega)$
as constructed in Theorem 1 is $\gamma$-continuous. As a consequence, $\Phi$ extends as a homomorphism in the Humilière-completion
$\widehat{\Phi}\colon\mathbb{R}\to\widehat{\text{Ham}}(M,\omega).$
Moreover, the map $\widehat{\Phi}$ is $1/\alpha$-periodic for an irrational number $\alpha\in\mathbb{R}$ and descends to a map, still denoted $\widehat{\Phi}$,
$\widehat{\Phi}\colon S^{1}\to\widehat{\text{Ham}}(M,\omega).$
This last map is injective and its image is closed for the $\gamma$-topology.
Proof.
As in the proof of Proposition 3.2, we have that
$\gamma(\Phi(r))=\gamma(S_{\alpha r}).$
Since the map $t\mapsto S_{\alpha t}$ is $1/\alpha$-periodic and $\widehat{\Phi}$ is a group homomorphism, we can conclude that $\widehat{\Phi}$ is also $1/\alpha$-periodic by the non-degeneracy of the $\gamma$-norm.
To show that the image is closed, we adapt the proof of Proposition 3.3 by noticing that the map $t\mapsto\gamma(S_{\alpha t})$ is $1/\alpha$-periodic, continuous and takes the value 0 only at integer multiples of $\frac{1}{\alpha}$. Hence, a sequence of maps $(\widehat{\Phi}(x_{n}))_{n}$ is a Cauchy sequence for the $\gamma$-topology implies that the sequence $(x_{n})_{n}$ is also Cauchy in $S^{1}$, we conclude the proof as in Proposition 3.3. ∎
We are still not able to answer Question 3 in this setting, but we have the following partial answer for $(\mathbb{C}P^{n},\omega_{FS})$ the standard symplectic projective space. We recall that in this case, the $\gamma$-norm is known to be continuous for the $C^{0}$-topology; Kawamoto has proved this fact in [Kaw22].
Proposition 3.5.
Let
$\widehat{\Phi}\colon S^{1}\to\widehat{\text{Ham}}(\mathbb{C}P^{n},\omega_{FS})$
be the extension defined in Proposition 3.4 and assume that the rotation number of $\Phi(1)$ is exponentially Liouville, then there exists an uncountable subgroup $G\subset S^{1}$ such that all the elements in $\widehat{\Phi}(G)$ are all homeomorphisms.
Question 4.
What kind of dynamical properties can the homeomorphisms of Proposition 3.5 have?
In the construction of the map $\Phi$ of Theorem 1, we say that the number $\alpha\in\mathbb{R}$ that is the limit of the sequence $(\alpha_{n})$ is the rotation number of the Hamiltonian diffeomorphism $\Phi(1)$. We also say that a real number $\alpha$ is exponentially Liouville if
$\forall c>0,\exists k,n\in\mathbb{Z}\text{ s.t. }0<|k\alpha-n|<e^{-ck}$
The set of exponentially Liouville numbers is residual and hence non-empty; see [GG18].
To prove Proposition 3.5, we need the following two lemmas and one corollary. We postpone their proof for the end of this section.
Lemma 3.6.
A complete metric space with no isolated points is uncountable.
Joksimovic and Seyfaddini have proved the following theorem in [JS24]. We recall first that a map $\varphi$ is said to be $C^{0}$-rigid if there exists a sequence of positive integers $n_{k}$ such that
$\varphi^{n_{k}}\xrightarrow[k\to+\infty]{C^{0}}Id.$
Theorem A (Joksimovic-Seyfaddini, [JS24]).
Let $(M,\omega)$ be a symplectic manifold with a Hamiltonian circle action $S$ and $\Phi$ the map constructed above. Assume that the rotation number of $\Phi(1)$ is exponentially Liouville. Then, the map $\Phi(1)$ is $C^{0}$-rigid.
This Theorem has the following implication.
Corollary 3.7.
Under the same assumption as in Theorem A, the set $\text{Im}(\Phi)\in\text{Ham}(M,\omega)$ has no isolated point for the $C^{0}$-topology.
Proof of Corollary 3.7.
From Theorem A, we get $(n_{k})_{k}$ a sequence of integers such that for the $C^{0}$-topology $\lim_{k\to\infty}\Phi(n_{k})=0$, this implies that $Id=\Phi(0)$ is not isolated, since $\Phi$ is a group homomorphism, then for every $q\in\mathbb{Q}$, the point $\Phi(q)$ is also not isolated. This finishes the proof of the corollary. ∎
We can now prove Proposition 3.5.
Proof of Proposition 3.5.
Since $\text{Im}(\Phi)$ is an abelian subgroup of $\text{Ham}(\mathbb{C}P^{n},\omega_{FS})$ and using the $C^{0}$-continuity of the $\gamma$-norm, its closure for the $C^{0}$-topology is also an abelian subgroup of $\widehat{\text{Ham}}(\mathbb{C}P^{n},\omega)$. We have the following inclusions
$\overline{\text{Im}(\Phi)}^{C^{0}}\subset\text{Im}(\widehat{\Phi})\subset\widehat{\text{Ham}}(\mathbb{C}P^{n},\omega).$
From Proposition 3.4, we get that
$G:=\widehat{\Phi}^{-1}\left(\overline{\text{Im}(\Phi)}^{C^{0}}\right)$
is a subgroup of $S^{1}$. From Corollary 3.7, we get that the $C^{0}$-closure of $\text{Im}(\Phi)$ has no isolated points, and from Lemma 3.6 it is uncountable. This means that $G$ is also uncountable, and since its image under $\widehat{\Phi}$ lies in $\text{Ham}(\mathbb{C}P^{n},\omega)$, the elements of the image are true homeomorphisms. ∎
If one can show the $C^{\infty}$-rigidity of maps obtained by the Anosov-Katok method with exponentially Liouville rotation number, one could extend Proposition 3.5 to additionally find an uncountable amount of diffeomorphisms in $\text{Im}(\widehat{\Phi})$.
We now prove Lemma 3.6.
Proof of Lemma 3.6.
Let $(X,d)$ be a complete metric space with no isolated points. We construct $2^{\omega}$ Cauchy sequences that converge to pairwise different points. We inductively construct sets $A_{n}$ of $2^{n}$ pairwise disjoint balls as follows. Let $B$ be a ball of radius $\varepsilon$. Since no isolated points exist, we find $B_{1},B_{2}\subset B$ two non-empty disjoint balls with a radius smaller than $\varepsilon\over 2$. We define
$A_{n+1}=\bigcup_{B\in A_{n}}\{B_{1},B_{2}\}$
while $A_{0}=\{B\}$ is a singleton of an arbitrary ball. For any decreasing sequence
$E_{0}\supset E_{2}\supset...\supset E_{n}\supset...$
of balls in $A=\bigcup_{n\in\mathbb{Z}_{\geq}0}A_{n}$ we have that $E=\bigcap_{n\in\mathbb{Z}_{\geq 0}}E_{n}$ is non-empty (since $X$ is complete). There are $2^{\omega}$ such sequences, which all lead to pairwise disjoint sets $E$. This finished the proof. ∎